Using definitions, check whether the following matrices are positive definite or positive semidefinite: 1 . 1 (2) 4-61. 8-6].a and b
Answer #1

A = lo clearly eigen-values of taxe = 11 Since both eigen value cite the and non Zedo thin A is tve definite Matrix B EigenE = Ⓡ . clearly buth eigen-valus are so semi-definite o.and.o © A = u 0 Mi =10 ľ il Ma = 0 6 My - M.. Me and My be first secoCe:? M, T = Ma - la in M? 1 C (1) + (0-1)=1-1=0 cleadly mi=iso. Ma also Mi=130, Maal so and Maco 80 c is semi-positive defini